OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2). -20 1.6 -1 0

OpenStudy (alexandervonhumboldt2):

substitute -2 for x and solve

OpenStudy (alexandervonhumboldt2):

\[h(-2)=-5*(-2)-10=?\]

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle h(x)=-5x-10 }\) If I, for instance, we to solve for h(-6), then this is what I would do... Step 1: Plug in x=-6 \(\color{#000000 }{ \displaystyle h(\color{#ff0000 }{-6})=-5\color{#ff0000 }{(-6)}-10 }\) Ste[ 2: Simplify this expression (We know that 6•5=30, and when the two numbers are negative, your product is still positive. So (-5)•(-6)=30, still) \(\color{#000000 }{ \displaystyle h(\color{#ff0000 }{-6})=30-10 }\) And your fuinal answer would be, \(\color{#000000 }{ \displaystyle h(\color{#ff0000 }{-6})=20}\)

OpenStudy (solomonzelman):

{{ That \(\Uparrow\) is just an EXAMPLE. }}